Four firefighters were discussing the number of fires they extinguished for the month of February. Rafael extinguished 25% fewer fires than Louis in February. William extinguished 50% more fires than Rafael. Albert extinguished 20% fewer fires than Rafael and William combined. If Louis extinguished 6 fires per week, how many fires did all four men extinguish during the entire month?
: Four firefighters were discussing the number of fires they extinguished for the month of February. Rafael extinguished 25% fewer fires than Louis in February. William extinguished 50% more fires than Rafael. Albert extinguished 20% fewer fires than Rafael and William combined. If Louis extinguished 6 fires per week, how many fires did all four men extinguish during the entire month?

Answer by ankor@dixie-net.com(4538)

Four firefighters were discussing the number of fires they extinguished for the month of February. Rafael extinguished 25% fewer fires than Louis in February. William extinguished 50% more fires than Rafael. Albert extinguished 20% fewer fires than Rafael and William combined. If Louis extinguished 6 fires per week, how many fires did all four men extinguish during the entire month?
:
Write an equation for each statmente (using decimals for %); simplify equations:
:
" Rafael extinguished 25% fewer fires than Louis in February."
R = L - .25L
R = .75L
:
" William extinguished 50% more fires than Rafael."
W = R + .5R
W = 1.5R
:
"Albert extinguished 20% fewer fires than Rafael and William combined."
A = (R+W) - .2(R+W)
A = .8(R+W)
:
"If Louis extinguished 6 fires per week,"
L = 4 * 6
L = 24 fires per month
:
how many fires did all four men extinguish during the entire month?
Substitute 24 for L
R = .75L
R = .75(24)
R = 18 fires for Raphael
and
W = 1.5R
Substitute 18 for R
W = 1.5 * 18
W = 27 fires for William
and
A = .8(R+W)
A = .8(18 + 27)
A = .8(45)
A = 36 fires for Albert
:
Four men total: 24 + 18 + 27 + 36 = 105 fires